OpenStudy (anonymous):

how to integrate arctan 3x.dx

OpenStudy (anonymous):

thanks ;)

myininaya (myininaya):

Here;s something else: (another way) Let 3x=tan(u) <- we need to make a triangle for this later =>3 dx=sec^2(u) du so we have \[\int\limits_{}^{}\arctan(3x) dx=\int\limits_{}^{}\arctan(\tan(u)) \frac{1}{3} \sec^2(u) du\] \[=\frac{1}{3}\int\limits_{}^{}u \sec^2(u) du\] so now we do integration by parts! \[=\frac{1}{3}(u \tan(u)-\int\limits_{}^{}\tan(u) du)+C\] \[=\frac{1}{3} u \tan(u)-\frac{1}{3}\int\limits_{}^{}\frac{\sin(u)}{\cos(u)} du +C\] we need a simple substitution here! :) let b=cos(u) => db=-sin(u) du \[=\frac{1}{3}u \tan(u)-\frac{1}{3}\int\limits_{}^{}\frac{-db}{b}+C\] \[=\frac{1}{3} u \tan(u)+\frac{1}{3} \ln|b|+C\] \[=\frac{1}{3}u \tan(u)+\frac{1}{3} \ln|\cos(u)|+C\] now we need this back in terms of x we call the first substitution so we have 3x=tan(u)  We can find the hyp by using the Pythagorean thm \[hyp=\sqrt{(3x)^2+1^2}=\sqrt{9x^2+1}\] \[=\frac{1}{3} \tan^{-1}(3x) \cdot 3x+\frac{1}{3} \ln|\frac{1}{\sqrt{9x^2+1}}|+C\] \[=x \tan^{-1}(3x)+\frac{1}{3} \ln|\frac{1}{\sqrt{9x^2+1}}|+C\]